# Robust $\mathcal{H}_{\infty}$ control and worst-case search in constrained parametric space

### v2

@@ III-A Problem formulation
As in [apkarian2015a], worst-case stability is studied by maximizing the spectral abscissa of the closed-loop matrix $A(\delta)$. The problem (4) can also be written as the following constrained optimization problem:
@@ III-B Properties of the objective functions
In this section, we consider a generic transfer function written as the Linear Fractional Transformation (LFT) [Zhou1996, Chapter 10] $T_{zw}(s,\delta)=\mathcal{F}_{u}(M(s),\Delta)$, as depicted in Figure 1, where $\Delta=\textrm{diag}(\delta_{i}I_{n_{i}})$ ($n_{i}$ is the number of repetitions of $\delta_{i}$). This representation will be used to characterize either the internal stability, or the $\mathcal{H}_{\infty}$ or $\mathcal{H}_{2}$ norm of the transfer from $w$ to $z$; practical problems e.g. (2) can also be addressed after an appropriate partitioning.
Figure 1: LFT representation of $T_{zw}$
@@
to have no direct transfer from $\tilde{p}$ to $\tilde{q}$. This change of variable is the same as in [Zhou1996, Section 12.3.4], with the uncertainty block instead of a feedback controller, and will be useful to study the $\mathcal{H}_{2}$ norm.
@@ III-B1 $\mathcal{H}_{2}$ norm (function $h_{2}$)
Proof.
The assumption $D_{22}=0$ is necessary to ensure finiteness of the $\mathcal{H}_{2}$ norm in the nominal case $\Delta=0$. The assumption $D_{12}=0$ or $D_{21}=0$ is done for the same reason for $\Delta\neq 0$. The differentiability of $h_{2}$ in $\mathcal{D}_{s}=\{\delta\in\mathcal{D},T_{zw}(s,\delta)\text{ is stable}\}$ was shown in [Rautert1997, Section 3]; it was originally derived for a controller rather than an uncertainty block, but the validity of the result extends straightforwardly to the latter. Let us write the state-space matrices of the transfer from $w$ to $z$ after closing the loop between $p$ and $q$:
@@
The derivative of $h_{2}$ at point $\delta$ for the variable $\widetilde{\Delta}$, noted $h_{2}^{\prime}(\widetilde{\Delta})d\widetilde{\Delta}$, is obtained from [apkarian2009, Lemma 3.1] or equivalently [Arzelier2010, Section 2.1], also originally derived for a controller and adapted here for the uncertain block $\widetilde{\Delta}$:
@@
and using the same notation as in [apkarian2015a]:
@@ III-B2 $\mathcal{H}_{\infty}$ norm (function $h_{\infty}$)
Proposition 5 (Regularity of $h_{\infty}$ [Boyd91, apkarian2006, apkarian2015a])
a) $h_{\infty}$ is locally Lipschitz and thus Clarke subdifferentiable (in the sense of [Clarke90]) everywhere in
@@
Proposition 5.a) is proved in [apkarian2015a, Proposition 1] and follows the work of [Boyd91] and [apkarian2006]. The general subdifferential expression is provided in these references, but we will not use it in this paper. b) is proved in [apkarian2015a, Proposition 2]. c) immediately follows from a), and the expression is also provided in [apkarian2015a, Remark 5]: using the same notations, let us define $T_{qw}$ and $T_{zp}$ as:
@@
with $\star$ denoting the redheffer star product. The transfer functions $T_{zw}$, $T_{qw}$ and $T_{zp}$ correspond to the transfer between the inputs $p,w$ and outputs $q,z$ of the matrix $M(s)$ as in Figure 1, but after closing the loop with the block $\Delta$. If $\omega_{0}$ is the single active frequency, such that the multiplicity of $\bar{\sigma}(T_{zw}(j\omega_{0},\delta))$ is 1, $h_{\infty}$ is differentiable and, as shown in [apkarian2015a], its derivatives read:
@@
where $\omega_{0}$ is an active frequency – but not necessarily the only one, and the multiplicity of $\bar{\sigma}(T_{zw}(j\omega_{0},\delta))$ is not necessarily 1 – and $p_{\omega_{0}}$ and $q_{\omega_{0}}$ are the corresponding normalized right and left singular vectors of $T_{zw}(j\omega_{0},\delta)$. Then, $d_{\delta}$ is a subgradient of $h_{\infty}$ (this is straightforward from the expression of the subgradient given in [apkarian2015a, Section IV-1]); by Proposition 2 and Theorem 3, it can legitimately be used in SQP as if it were the gradient of a differentiable function.
@@
In contrast, minimizing the lower-$C^{1}$ function $h_{\infty}(\delta)$, i.e. minimizing the $\mathcal{H}_{\infty}$ norm, would be a more complex, genuinely nonsmooth problem. In particular, an unconstrained minimum may occur at a point where many subgradients are not zero (as a more simple example, consider $f:x\rightarrow|x|$, which is also lower-$C^{1}$). For this problem, nonsmooth techniques in the literature include multidirectionnal search [apkarian2006MDS], gradient sampling [Burke2005], or Clarke subdifferential-based [apkarian2006, Bompart2007a]. This is also pointed out in [apkarian2015a] as a fundamental difference between min-max and min-min (or max-max) programs, the former (e.g. minimizing the $\mathcal{H}_{\infty}$ norm) being genuinely non-smooth while the latter (e.g. maximizing the $\mathcal{H}_{\infty}$ norm or equivalently minimizing $h_{\infty}$) having the favorable properties demonstrated above.
@@ III-B3 Spectral abscissa (function $a$)
The case of the spectral abscissa is slightly less favorable than the $\mathcal{H}_{\infty}$ norm. We recall that simple eigenvalues (i.e. those of multiplicity one) are differentiable [Overton1988]. As in [apkarian2015a], we call ”active eigenvalues” all eigenvalues whose real part is equal to the spectral abscissa.
Proposition 6 (Regularity of $a$ [apkarian2015a, Bompart2007])
a) If all active eigenvalues of $A(\delta)$ are semi-simple (i.e. their algebraic and geometric multiplicity are equal), then $a$ is locally Lipschitz and thus Clarke subdifferentiable at $\delta$.
@@
Proposition 6.a) was established in [apkarian2015a] following the earlier work of [Bompart2007]. These references build on the study of [Burke94] concerning the directional derivatives of the spectral abscissa. b) is also shown in [apkarian2015a], it simply follows from the fact that the spectral abscissa is the maximum of the real parts of the eigenvalues, which are differentiable when the eigenvalues are simple [Overton1988]. For c), the expression (20) is straightforward from the expression of the Clarke subdifferential given in [apkarian2015a].
@@
Additionally, if there are two active eigenvalues and they are complex conjugate eigenvalues, $a$ is also differentiable and the gradient given above remains valid for either of the two eigenvalues [Bompart2007].
@@
Although the assumption of semi-simplicity is required for the definition of the Clarke subdifferential, the numerical experiments reported in [apkarian2015a] show that their bundle approach, based on Clarke subdifferentials, performs well in practice even without explicitly verifying this assumption.
@@
As for the $\mathcal{H}_{\infty}$ norm, minimizing the spectral abscissa is a more complex problem; in the literature, this is addressed with bundle methods [Oustry2000, Noll2005], gradient bundle [Burke2002], gradient sampling [Burke2005], Clarke subdifferential-based [Bompart2007].
@@
III-C Choice of optimization algorithms
Problems (12), (13) and (14) are constrained optimization problems involving the nonsmooth functions $a$, $h_{\infty}$, and the differentiable function $h_{2}$. We propose to use either Particle Swarm Optimization or Monte-Carlo sampling for global exploration of the parametric space, and then Sequential Quadratic Programming for local exploitation. As demonstrated in Section II, this procedure will converge to KKT points of the optimization problems.
@@ III-C1 Sequential Quadratic Programming (SQP)
SPQ is described in Algorithm 1 and, as discussed in Section II, may be used to minimize upper-$C^{1}$ functions. In our numerical application, we used MATLAB routine fmincon [matlaboptim]. In fmincon, the matrix $H$ of the quadratic problem (10) serves as an approximation of the Hessian matrix of the Lagrangian, iteratively updated with the Broyden-Fletcher-Goldfarb-Shanno (BFGS) method. It is worth noting that BFGS was observed to perform well in [Bompart2007] even for a non-smooth objective function (the spectral abscissa).
@@ III-C2 Particle swarm optimization (PSO)
and problems (12) and (14) can be treated similarly. The parameter $\tau$ can be set to a high value, and iteratively increased if necessary until a solution $\delta^{*}$ verifying all constraints $c_{i}(\delta^{*})\leq 0$ is found. The bounds on the $\delta_{i}$ are enforced by projecting the particles positions into the admissible space at each iteration. In our numerical testings, we used MATLAB routine pso described in [matlabpso]; PSO was used as a method to explore the parametric space (and the choice of parameters was done accordingly, as discussed in Section V-C), with local exploitation being performed by a subsequent SQP.
@@ III-C3 Monte-Carlo (MC) sampling
where $\gamma$ and $\epsilon$ serve as confidence and accuracy levels to guarantee the probability [Tempo1996]:
@@
for an objective function $f$ (in our case: $a(\delta)$, $h_{\infty}(\delta)$ or $h_{2}(\delta)$). Note that this bound is independent from the number of uncertain parameters, from their variation ranges or distributions.
@@ IV-A Robust controller optimization algorithm
The Algorithm 2, described in a form that solves the standard $\mathcal{H}_{\infty}$ control problem (1) – the multi-objectives formulation is tackled in Section IV-B – follows the approach proposed in [apkarian2015a] that relies on iteratively constructing a subset $\mathcal{D}_{a}\subset\mathcal{D}$ of so-called active configurations as follows. Step 1 optimizes the controller for all current active configurations. It uses the nonsmooth optimization algorithm of [apkarian2006] implemented in the routine systune. Step 2 addresses the worst-case search in stability as described in Section III; if a worst-case configuration, i.e. a Karush-Kuhn-Tucker (KKT) point of the optimization problem, is found that leads to a spectral abscissa greater than a tolerance $\alpha_{max}$, it is added to $\mathcal{D}_{a}$. It is also possible to identify multiple KKT points during Step 2. Similarly, Step 3 tackles the worst-case search in performance and adds to $\mathcal{D}_{a}$ the worst-case configuration(s) that degrade the performance by more than a tolerance $\epsilon$. Finally, Step 4 performs a more exhaustive worst-case search, meaning that the optimization parameters are tuned to increase coverage of the parameter space and ensure a more thorough exploration of potential worst-case configurations, while in Steps 2 and 3 they are chosen to prioritize fast convergence to KKT points; additional quantitative details are provided in the application of Section V-D. For Steps 2, 3 and 4, the global exploration may be performed either with Particle Swarm Optimization (PSO), or Monte-Carlo (MC) if stochastic guarantees are sought; in both cases, a subsequent Sequential Quadratic Programming (SQP) refines the search until finding a KKT point.
@@ IV-B Multi-objective formulation
since it is equivalent to a single objective problem (e.g. as in [Apkarian2006b, apkarian2017]):
@@
More generally, we consider the multi-objectives problem (2). As explained in [apkarian2014], the controller optimization of Step 1 treats it as :
@@ V-A Plant model and uncertain parameters
The benchmark used in this study is designed to be challenging regarding robustness. We consider a spacecraft composed of one rigid body connected to two large flexible solar panels. The clamped-free natural frequencies are 0.20, 0.50, 0.75 and 0.85 rad/s for each panel. The resulting 8 nominal free-free frequencies are between 0.20 and 4.2 rad/s for the whole spacecraft. Uncertainties are considered on various mechanical properties:
@@
The spacecraft model $G(s,\delta)$ is built with the multibody approach detailed in [alazard2008linear] and implemented in the Satellite Dynamics Toolbox Library (SDTlib) [Alazard2020, Sanfedino2023sdt]. The block $\Delta_{G}$ of the Linear Fractional Representation (LFR) model, such that $G(s,\delta)=\mathcal{F}_{u}(M_{G}(s),\Delta_{G})$ (upper LFT), has size 204. We emphasize that this model contains all parametric configurations in a continuous manner, defined at each subsystem level based on the physics equations without any interpolation or other approximation (besides the linearity assumption). The LFR model of the spacecraft includes all configurations for $1\leq\delta_{i}\leq 1$ even though only $\sim 22\%$ of them respect the nonlinear constraint (25).

### v4

@@ III-A Problem formulation
As in [3], worst-case stability is studied by maximizing the spectral abscissa of the closed-loop matrix $A(\delta)$. The problem (4) can also be written as the following constrained optimization problem:
@@ III-B Properties of the objective functions
In this section, we consider a generic transfer function written as the Linear Fractional Transformation (LFT) [54, Chapter 10] $T_{zw}(s,\delta)=\mathcal{F}_{u}(M(s),\Delta)$, as depicted in Figure 1, where $\Delta=\textrm{diag}(\delta_{i}I_{n_{i}})$ ($n_{i}$ is the number of repetitions of $\delta_{i}$). This representation will be used to characterize either the internal stability, or the $\mathcal{H}_{\infty}$ or $\mathcal{H}_{2}$ norm of the transfer from $w$ to $z$; practical problems e.g. (2) can also be addressed after an appropriate partitioning.
Figure 1: LFT representation of $T_{zw}(s,\delta)$
@@
to have no direct transfer from $\tilde{p}$ to $\tilde{q}$. This change of variable is the same as in [54, Section 12.3.4], with the uncertainty block instead of a feedback controller, and will be useful to study the $\mathcal{H}_{2}$ norm.
@@ III-B1 $\mathcal{H}_{2}$ norm (function $h_{2}$)
While the proof mainly follows previously established results cited thereafter, it is necessary to present the following equations to obtain the explicit gradient expression (18). The assumption $D_{22}=0$ is necessary to ensure finiteness of the $\mathcal{H}_{2}$ norm in the nominal case $\Delta=0$. The assumption $D_{12}=0$ or $D_{21}=0$ is done for the same reason for $\Delta\neq 0$. The differentiability of $h_{2}$ in $\mathcal{D}_{s}=\{\delta\in\mathcal{D},T_{zw}(s,\delta)\text{ is stable}\}$ was shown in [45, Section 3]; it was originally derived for a controller rather than an uncertainty block, but the validity of the result extends straightforwardly to the latter. Let us write the state-space matrices of the transfer from $w$ to $z$ after closing the loop between $p$ and $q$:
@@
The derivative of $h_{2}$ at point $\delta$ for the variable $\widetilde{\Delta}$, noted $h_{2}^{\prime}(\widetilde{\Delta})d\widetilde{\Delta}$, is obtained from [6, Lemma 3.1] or equivalently [12, Section 2.1], also originally derived for a controller and adapted here for the uncertain block $\widetilde{\Delta}$:
@@
and using the same notation as in [3]:
@@ III-B2 $\mathcal{H}_{\infty}$ norm (function $h_{\infty}$)
Proposition 5 (Regularity of $h_{\infty}$ [16, 8, 3])
a) $h_{\infty}$ is locally Lipschitz and Clarke subdifferentiable (in the sense of [22]) everywhere in
@@
Proposition 5.a) is proved in [3, Proposition 1] and follows the work of [16] and [8]. The general subdifferential expression is provided in these references, but we will not use it in this paper. b) is proved in [3, Proposition 2]. c) immediately follows from a), and the expression is also provided in [3, Remark 5]: using the same notations, let us define $T_{qw}$ and $T_{zp}$ as:
@@
with $\star$ denoting the redheffer star product. The transfer functions $T_{zw}$, $T_{qw}$ and $T_{zp}$ correspond to the transfer between the inputs $p,w$ and outputs $q,z$ of the matrix $M(s)$ as in Figure 1, but after closing the loop with the block $\Delta$. If $\omega_{0}$ is the single active frequency, such that the multiplicity of $\bar{\sigma}(T_{zw}(j\omega_{0},\delta))$ is 1, $h_{\infty}$ is differentiable and, as shown in [3], its derivatives read:
@@
where $\omega_{0}$ is an active frequency – but not necessarily the only one, and the multiplicity of $\bar{\sigma}(T_{zw}(j\omega_{0},\delta))$ is not necessarily 1 – and $p_{\omega_{0}}$ and $q_{\omega_{0}}$ are the corresponding normalized right and left singular vectors of $T_{zw}(j\omega_{0},\delta)$. Then, $d_{\delta}$ is a subgradient of $h_{\infty}$ (this is straightforward from the expression of the subdifferential given in [3, Section IV-1]); by Propositions 2 and 3, it can legitimately be used in SQP as if it were the gradient of a differentiable function.
@@
In contrast, minimizing the lower-$C^{1}$ function $h_{\infty}(\delta)$, i.e. minimizing the $\mathcal{H}_{\infty}$ norm, would be a more complex, genuinely nonsmooth problem. In particular, an unconstrained minimum may occur at a point where many subgradients are not zero (as a more simple example, consider $f:x\rightarrow|x|$, which is also lower-$C^{1}$). For this problem, nonsmooth techniques in the literature include multidirectionnal search [7], gradient sampling [18], or Clarke subdifferential-based [8, 15]. This is also pointed out in [3] as a fundamental difference between min-max and min-min (or max-max) programs, the former (e.g. minimizing the $\mathcal{H}_{\infty}$ norm) being genuinely non-smooth while the latter (e.g. maximizing the $\mathcal{H}_{\infty}$ norm or equivalently minimizing $h_{\infty}$) having the favorable properties demonstrated above.
@@ III-B3 Spectral abscissa (function $a$)
The case of the spectral abscissa is slightly less favorable than the $\mathcal{H}_{\infty}$ norm. We recall that simple eigenvalues (i.e. those of multiplicity one) are differentiable [41]. As in [3], we call ”active eigenvalues” all eigenvalues whose real part is equal to the spectral abscissa.
Proposition 6 (Regularity of $a$ [3, 14])
a) If all active eigenvalues of $A(\delta)$ are semi-simple (i.e. their algebraic and geometric multiplicity are equal), then $a$ is locally Lipschitz and Clarke subdifferentiable at $\delta$.
@@
Proposition 6.a) was established in [3] following the earlier work of [14]. These references build on the study of [19] concerning the directional derivatives of the spectral abscissa. b) is also shown in [3], it simply follows from the fact that the spectral abscissa is the maximum of the real parts of the eigenvalues, which are differentiable when the eigenvalues are simple [41]. For c), the expression (20) is straightforward from the expression of the Clarke subdifferential given in [3].
@@
Additionally, if there are two active eigenvalues and they are complex conjugate eigenvalues, $a$ is also differentiable and the gradient given above remains valid for either of the two eigenvalues [14].
@@
Although the assumption of semi-simplicity is required for the definition of the Clarke subdifferential, the numerical experiments reported in [3] show that their bundle approach, based on Clarke subdifferentials, performs well in practice even without explicitly verifying this assumption.
@@
As for the $\mathcal{H}_{\infty}$ norm, minimizing the spectral abscissa is a more complex problem; in the literature, this is addressed with bundle methods [40, 39], gradient bundle [17], gradient sampling [18], Clarke subdifferential-based [14].
@@
III-C Global exploration, local exploitation
Problems (12), (13) and (14) are constrained optimization problems involving the nonsmooth functions $a$ and $h_{\infty}$, and the differentiable function $h_{2}$. As demonstrated in Section II, Sequential Quadratic Programming may be used, as it will converge to KKT points of the optimization problems under the upper-$C^{1}$ assumption. This constitutes the main theoretical contribution of the present paper. From a practical perspective, however, we additionally investigate the use of Particle Swarm Optimization or Monte-Carlo sampling to perform a global exploration of the parameter space, with SQP subsequently ensuring efficient local exploitation.
@@ III-C1 Sequential Quadratic Programming (SQP)
SPQ is described in Algorithm 1 and, as discussed in Section II, may be used to minimize upper-$C^{1}$ functions. In our numerical application, we used MATLAB routine fmincon [23]. In fmincon, the matrix $H$ of the quadratic problem (10) serves as an approximation of the Hessian matrix of the Lagrangian, iteratively updated with the Broyden-Fletcher-Goldfarb-Shanno (BFGS) method. It is worth noting that BFGS was observed to perform well in [14] even for a non-smooth objective function (the spectral abscissa).
@@ III-C2 Particle swarm optimization (PSO)
and problems (12) and (14) can be treated similarly. The parameter $\tau$ can be set to a high value, and iteratively increased if necessary until a solution $\delta^{*}$ verifying all constraints $c_{i}(\delta^{*})\leq 0$ is found. The bounds on the $\delta_{i}$ are enforced by projecting the particles positions into the admissible space at each iteration. In our numerical testings, we used MATLAB routine pso described in [42]; PSO was used as a method to explore the parametric space (and the choice of parameters was done accordingly, as discussed in Section V-C), with local exploitation being performed by a subsequent SQP.
@@ III-C3 Monte-Carlo (MC) sampling
where $\gamma$ and $\epsilon$ serve as confidence and accuracy levels to guarantee the probability [53]:
@@
for an objective function $f$ (in our case: $a(\delta)$, $h_{\infty}(\delta)$ or $h_{2}(\delta)$). Note that this bound is independent from the number of uncertain parameters, from their variation ranges or distributions. Since this paper addresses deterministic worst-case search, it is natural to draw parameters from uniform distributions; note, however, that removing some inadmissible configurations implicitly modifies these distributions.
@@ IV-A Robust controller optimization algorithm
The Algorithm 2, described in a form that solves the standard $\mathcal{H}_{\infty}$ control problem (1) – the multi-objectives formulation is tackled in Section IV-B – follows the approach proposed in [3] that relies on iteratively constructing a subset $\mathcal{D}_{a}\subset\mathcal{D}$ of so-called active configurations as follows.
Step 1 optimizes the controller for all current active configurations. It uses the nonsmooth optimization algorithm of [8] implemented in the routine systune. Step 2 addresses the worst-case search in stability as described in Section III; if a worst-case configuration, i.e. a Karush-Kuhn-Tucker (KKT) point of the optimization problem, is found that leads to a spectral abscissa greater than a tolerance $\alpha_{max}$, it is added to $\mathcal{D}_{a}$. It is also possible to identify multiple KKT points during Step 2. Similarly, Step 3 tackles the worst-case search in performance and adds to $\mathcal{D}_{a}$ the worst-case configuration(s) that degrade the performance by more than a tolerance $\epsilon$. Finally, Step 4 performs a more exhaustive worst-case search, meaning that the optimization parameters are tuned to increase coverage of the parameter space and ensure a more thorough exploration of potential worst-case configurations, while in Steps 2 and 3 they are chosen to prioritize fast convergence to KKT points; additional quantitative details are provided in the application of Section V-D. For Steps 2, 3 and 4, the global exploration may be performed either with Particle Swarm Optimization (PSO), or Monte-Carlo (MC) if stochastic guarantees are sought; in both cases, a subsequent Sequential Quadratic Programming (SQP) refines the search until finding a KKT point.
@@ IV-B Multi-objective formulation
since it is equivalent to a single objective problem (e.g. as in [9, 10]):
@@
More generally, we consider the multi-objectives problem (2). As explained in [4], the controller optimization of Step 1 treats it as :
@@ V-A Plant model and uncertain parameters
The benchmark used in this study is designed to be challenging regarding robustness. We consider a spacecraft composed of one rigid body connected to two large flexible solar panels. The clamped-free natural frequencies are between 0.20 and 0.85 rad/s for each panel. The resulting 8 nominal free-free frequencies are between 0.20 and 4.2 rad/s for the whole spacecraft. Uncertainties are considered on various mechanical properties:
@@
The spacecraft model $G(s,\delta)$ is built with the multibody approach detailed in [1] and implemented in the Satellite Dynamics Toolbox Library (SDTlib) [2, 48]. The block $\Delta_{G}$ of the Linear Fractional Representation (LFR) model, such that $G(s,\delta)=\mathcal{F}_{u}(M_{G}(s),\Delta_{G})$ (upper LFT), has size 204. We emphasize that this model contains all parametric configurations in a continuous manner, defined at each subsystem level based on the physics equations without any interpolation or other approximation (besides the linearity assumption). The LFR model of the spacecraft includes all configurations for $1\leq\delta_{i}\leq 1$ even though only $\sim 22\%$ of them respect the nonlinear constraint (25).
Remark 8.
The constraint (25) is not differentiable everywhere, which contradicts the assumptions of Section 2. However, since this is the formulation that naturally arises in our application, we still test it numerically.